OpenStudy (alexismiichaela):

I need help with figuring out angles using Pythagorean theorem. For example, being given a triangle, with all three side's measurements given, and being asked to solve for the missing angles. (No angle's are given)

OpenStudy (3mar):

May I help?

OpenStudy (alexismiichaela):

Yes please!

OpenStudy (3mar):

With my pleasure! What did you get?

OpenStudy (alexismiichaela):

Like what is the question?

OpenStudy (3mar):

do you have it or what?

OpenStudy (alexismiichaela):

Yeah sorry, The triangles sides are 16km, 19km, and 19km again. And I need to calculate the angles.

OpenStudy (3mar):

But that triangle is not a right one!!

OpenStudy (alexismiichaela):

So then it wouldn't be Pythagorean theorum? My options were that, Sine law, Cosine law, and SOHCAHTOA

OpenStudy (3mar):

So we can use the fist two not the third!

OpenStudy (alexismiichaela):

How does that work? I thought you needed at least one angle to use those?

OpenStudy (3mar):

Not a conditon! you can use Cosine law to get the first angle, then Sine law to get the second, the (sum=180) to get the third! Got that!

OpenStudy (alexismiichaela):

Ohh okay that makes sense!

OpenStudy (3mar):

Definitely!

OpenStudy (3mar):

Can you share your works/steps, please?

OpenStudy (alexismiichaela):

I just need a minute or two to see if I got this first!

OpenStudy (3mar):

Take your time.

OpenStudy (alexismiichaela):

I'm a little rusty with Cosine law as well, so just to make sure I'm setting this up right.. \[a ^{2} = b ^{2} + c ^{2} - 2 (b) (c) cosA \] is the equation, so if I were to plug in the numbers it would be \[19^{2} = 16^{2} + 19^{2} -2 (16) (19) cosA\]

OpenStudy (alexismiichaela):

is this right so far?

OpenStudy (3mar):

Let me see

OpenStudy (3mar):

\[\Huge\color{darkred}\checkmark\] so A=?? degree

OpenStudy (alexismiichaela):

yess

OpenStudy (3mar):

Excellent! you got A=##??

OpenStudy (alexismiichaela):

I think I did something wrong, I ended up with 361 = 9cosA

OpenStudy (alexismiichaela):

I'm not sure what to do

OpenStudy (3mar):

For me, I get almost A=65 Check again and I am with you

OpenStudy (alexismiichaela):

Can I send you my steps and you tell me what I did wrong?

OpenStudy (3mar):

Thank you for the medal!

OpenStudy (3mar):

Of course.

OpenStudy (alexismiichaela):

\[19^{2}=16^{2}+19^{2}-2(16)(19)cosA\] \[19^{2}=16^{2}+19^{2}-2(304)cosA\] \[361=256+361-2(304)cosA\] \[361=256+361-608cosA\] \[361=9cosA\]

OpenStudy (3mar):

Brilliant! Ma Sha' Allah

OpenStudy (alexismiichaela):

Thats all right??

OpenStudy (3mar):

but just last one! \[361=256+361-608cosA~~~~~~~~~~~~~~~-256=-608~\cos(A)\] \[\Huge \color{LightSalmon }{ \cos(A)=\frac{ -256 }{ -608 } =\frac{ 8 }{ 19 }=0.4211}\] \[\Huge \color{Tomato }{ A=65}\]

OpenStudy (3mar):

@alexismiichaela got it?

OpenStudy (alexismiichaela):

That's actually confusing me a little bit.. where did the two 361 values go?

OpenStudy (3mar):

cancel each other what if you subtract 361 from each side????

OpenStudy (alexismiichaela):

OH right, okay I'll take another look to make sure I absolutely understand it

OpenStudy (3mar):

Take your look

OpenStudy (alexismiichaela):

Okay so I get that, and I also understand how the 256 was transferred onto the right side of the equal sign, but how does Cos(A) get transferred to the right, and the 256 value, back to the left side

OpenStudy (3mar):

\[361=256+361-608cos(A)\] \[361-361=256+361-361-608cos(A)\] \[0=256-608cos(A)\] \[0+608cos(A)=256-608cos(A)+-608cos(A)\] \[608cos(A)=256\] \[\Large \cos(A)=\frac{ 256 }{ 608 }\] got that?

OpenStudy (alexismiichaela):

Yes! That looks better

OpenStudy (triciaal):